# Fairness-Aware Fine-Tuning of Vision-Language Models for Medical Glaucoma Diagnosis

### v2

@@ 1 Introduction
Glaucoma, affecting over 70 million people worldwide, is a leading cause of irreversible blindness tham2014glaucoma. Vision-language models (VLMs) now achieve expert-level performance on glaucoma diagnosis tasks including disease classification, visual question answering, and report generation llava_med, qwen2vl. However, a critical challenge remains: performance disparities—systematic differences in diagnostic accuracy, sensitivity, or specificity across demographic groups (race, ethnicity, gender). Obermeyer et al. obermeyer2019dissecting found that Optum’s risk prediction algorithm, deployed across major U.S. hospitals, assigned systematically lower scores to Black patients than equally-sick White patients, halving their access to care. Similarly, dermatology AI shows 10-20% lower accuracy on darker skin, delaying melanoma diagnosis in underserved populations daneshjou2022disparities. For glaucoma—where Hispanic populations face 2-3 higher prevalence—a 10% accuracy gap translates to hundreds of preventable blindness cases annually.
Prior fairness methods for medical imaging—data augmentation xu2023fairness, adversarial debiasing glocker2023algorithmic, loss regularization—target CNN architectures, leaving two critical research gaps. First, whether these interventions transfer to billion-parameter VLMs with multimodal reasoning remains unexplored. Second, current approaches require full fine-tuning of all parameters, which is computationally prohibitive and prone to overfitting on limited medical datasets (). While parameter-efficient methods like LoRA hu2022lora enable efficient adaptation, their integration with fairness objectives in medical VLMs is unaddressed. Moreover, standard metrics (e.g., Equalized Odds hardt2016equality) optimize for equal error rates but permit large accuracy gaps: a model satisfying Equalized Odds may have 70% accuracy for one group and 90% for another, violating the clinical expectation of equal diagnostic accuracy for all patients.
@@ 2.1 Medical Vision-Language Models
Recent advances in VLMs have enabled their application to medical imaging tasks. Specialized medical VLMs such as LLaVA-Med llava_med and Med-Flamingo moor2023medflamingo are trained on biomedical image-text pairs. LLaVA-Med achieved expert-level performance on medical VQA by fine-tuning LLaVA on curated medical datasets, while Med-Flamingo demonstrated few-shot learning capabilities through in-context examples. General-purpose VLMs like Qwen2.5-VL qwen2vl, LLaVA-v1.6 liu2023improvedllava, and GPT-4V are increasingly adapted to medical tasks through fine-tuning. Qwen2.5-VL’s dynamic resolution processing makes it particularly suitable for fine-grained medical image analysis. Despite strong performance on standard benchmarks, systematic fairness evaluation across demographic groups remains absent in existing VLM research. Our work addresses this gap by fine-tuning a general-purpose VLM (Qwen2.5-VL-7B) for glaucoma diagnosis and conducting comprehensive fairness evaluation across gender, race, and ethnicity attributes.
@@ 2.2 Fairness in Medical Image Analysis
Prior work on fairness in medical AI spans multiple approaches. Data-level methods include generative models for synthetic minority samples xu2023fairness and balanced sampling strategies. However, synthetic data may introduce distributional shift and fail to capture real-world minority group characteristics. Loss-level methods include disentanglement learning locatello2019fairness, mutual information regularization, and adversarial debiasing glocker2023algorithmic. FairSeg addresses fairness in medical image segmentation using error-bound scaling to reweight pixel losses. These methods primarily target CNN-based architectures and require full model fine-tuning (100% parameters trainable), limiting their applicability to billion-parameter VLMs on limited medical datasets. In contrast, our LoRA-based approach fine-tunes only 0.06% of parameters, enabling fairness optimization without overfitting.
@@ 2.3 Parameter-Efficient Fine-Tuning for Fairness
LoRA hu2022lora enables efficient adaptation of large pre-trained models through low-rank weight updates where , with . FairTune fairtune2024 uses bi-level optimization to select which parameters to fine-tune for fairness, achieving 20-30% parameter reduction but still requiring millions of trainable parameters. FairLoRA fairlora2024 introduces loss variance regularization () to reduce group-wise loss disparities, but focuses on vision-only models (ResNet, ViT) and Equalized Odds rather than accuracy parity.
@@ 2.4 Fairness Metrics in Healthcare AI
Common fairness metrics include: Demographic Parity (, equal positive prediction rates), which is unsuitable for medical diagnosis as disease prevalence varies across demographics; Equalized Odds hardt2016equality (, , equal true/false positive rates), which ensures equal error rates but permits large accuracy gaps (e.g., Acc, Acc can satisfy Equalized Odds if both groups have TPR=0.8, FPR=0.2); and Equal Opportunity (equal TPR only), which ignores false positives.
@@ 3.1 Problem Formulation
Why MaxAccGap? Unlike distributional fairness metrics such as Equalized Odds hardt2016equality, MaxAccGap offers direct clinical interpretability: physicians and patients intuitively understand equal diagnostic accuracy, and MaxAccGap directly translates to clinical impact through the number of misdiagnoses in each demographic group.
@@ 3.2 LoRA for Vision-Language Models
For a pre-trained weight matrix in the VLM’s attention layers, LoRA hu2022lora approximates the weight update as:
@@
Implementation. Following standard practice for decoder-only VLMs brown2020gpt3, touvron2023llama, we use the final token’s hidden state for binary classification, as it has attended to all preceding tokens (vision features + text prompt) under causal attention.
@@ 4.1 Dataset and Setup
Dataset. We use the Harvard Glaucoma Fairness Dataset tham2014glaucoma containing 10,000 fundus images with demographic annotations across three sensitive attributes: gender (Male 42.9%, Female 57.1%), race (White 76.9%, Black 14.9%, Asian 8.2%), and ethnicity (Non-Hispanic 90.3%, Hispanic 4.3%, Unknown 5.4%). Each image is paired with a binary glaucoma diagnosis label and a clinical question (“Does the patient have glaucoma?”). The dataset is split into 7,000 training, 1,000 validation, and 2,000 test samples, maintaining demographic distributions across splits.
@@
Model. We fine-tune Qwen2.5-VL-7B-Instruct qwen2vl (8.3B parameters) with LoRA (, , dropout=) applied to query, key, value, and output projection layers () across all 32 attention layers. This configuration yields 20M trainable parameters (0.24% of the full model), striking a balance between expressiveness and parameter efficiency for fairness optimization on our medical dataset ( training samples). We freeze the vision encoder to preserve pre-trained visual features while adapting the language model for medical diagnosis.

### v3

@@ 1 Introduction
Glaucoma, affecting over 70 million people worldwide, is a leading cause of irreversible blindness 13. Vision-language models (VLMs) now achieve expert-level performance on glaucoma diagnosis tasks including disease classification, visual question answering, and report generation 7, 15. However, a critical challenge remains: performance disparities—systematic differences in diagnostic accuracy, sensitivity, or specificity across demographic groups (race, ethnicity, gender). Obermeyer et al. 11 found that Optum’s risk prediction algorithm, deployed across major U.S. hospitals, assigned systematically lower scores to Black patients than equally-sick White patients, halving their access to care. Similarly, dermatology AI shows 10-20% lower accuracy on darker skin, delaying melanoma diagnosis in underserved populations 2. For glaucoma—where Hispanic populations face 2-3 higher prevalence—a 10% accuracy gap translates to hundreds of preventable blindness cases annually.
Prior fairness methods for medical imaging—data augmentation 16, adversarial debiasing 4, loss regularization—target CNN architectures, leaving two critical research gaps. First, whether these interventions transfer to billion-parameter VLMs with multimodal reasoning remains unexplored. Second, current approaches require full fine-tuning of all parameters, which is computationally prohibitive and prone to overfitting on limited medical datasets (). While parameter-efficient methods like LoRA 6 enable efficient adaptation, their integration with fairness objectives in medical VLMs is unaddressed. Moreover, standard metrics (e.g., Equalized Odds 5) optimize for equal error rates but permit large accuracy gaps: a model satisfying Equalized Odds may have 70% accuracy for one group and 90% for another, violating the clinical expectation of equal diagnostic accuracy for all patients.
@@ 2.1 Medical Vision-Language Models
Recent advances in VLMs have enabled their application to medical imaging tasks. Specialized medical VLMs such as LLaVA-Med 7 and Med-Flamingo 10 are trained on biomedical image-text pairs. LLaVA-Med achieved expert-level performance on medical VQA by fine-tuning LLaVA on curated medical datasets, while Med-Flamingo demonstrated few-shot learning capabilities through in-context examples. General-purpose VLMs like Qwen2.5-VL 15, LLaVA-v1.6 8, and GPT-4V are increasingly adapted to medical tasks through fine-tuning. Qwen2.5-VL’s dynamic resolution processing makes it particularly suitable for fine-grained medical image analysis. Despite strong performance on standard benchmarks, systematic fairness evaluation across demographic groups remains absent in existing VLM research. Our work addresses this gap by fine-tuning a general-purpose VLM (Qwen2.5-VL-7B) for glaucoma diagnosis and conducting comprehensive fairness evaluation across gender, race, and ethnicity attributes.
@@ 2.2 Fairness in Medical Image Analysis
Prior work on fairness in medical AI spans multiple approaches. Data-level methods include generative models for synthetic minority samples 16 and balanced sampling strategies. However, synthetic data may introduce distributional shift and fail to capture real-world minority group characteristics. Loss-level methods include disentanglement learning 9, mutual information regularization, and adversarial debiasing 4. FairSeg addresses fairness in medical image segmentation using error-bound scaling to reweight pixel losses. These methods primarily target CNN-based architectures and require full model fine-tuning (100% parameters trainable), limiting their applicability to billion-parameter VLMs on limited medical datasets. In contrast, our LoRA-based approach fine-tunes only 0.06% of parameters, enabling fairness optimization without overfitting.
@@ 2.3 Parameter-Efficient Fine-Tuning for Fairness
LoRA 6 enables efficient adaptation of large pre-trained models through low-rank weight updates where , with . FairTune 3 uses bi-level optimization to select which parameters to fine-tune for fairness, achieving 20-30% parameter reduction but still requiring millions of trainable parameters. FairLoRA 12 introduces loss variance regularization () to reduce group-wise loss disparities, but focuses on vision-only models (ResNet, ViT) and Equalized Odds rather than accuracy parity.
@@ 2.4 Fairness Metrics in Healthcare AI
Common fairness metrics include: Demographic Parity (, equal positive prediction rates), which is unsuitable for medical diagnosis as disease prevalence varies across demographics; Equalized Odds 5 (, , equal true/false positive rates), which ensures equal error rates but permits large accuracy gaps (e.g., Acc, Acc can satisfy Equalized Odds if both groups have TPR=0.8, FPR=0.2); and Equal Opportunity (equal TPR only), which ignores false positives.
@@ 3.1 Problem Formulation
Why MaxAccGap? Unlike distributional fairness metrics such as Equalized Odds 5, MaxAccGap offers direct clinical interpretability: physicians and patients intuitively understand equal diagnostic accuracy, and MaxAccGap directly translates to clinical impact through the number of misdiagnoses in each demographic group.
@@ 3.2 LoRA for Vision-Language Models
For a pre-trained weight matrix in the VLM’s attention layers, LoRA 6 approximates the weight update as:
@@ 4.1 Dataset and Setup
Dataset. We use the Harvard Glaucoma Fairness Dataset 13 containing 10,000 fundus images with demographic annotations across three sensitive attributes: gender (Male 42.9%, Female 57.1%), race (White 76.9%, Black 14.9%, Asian 8.2%), and ethnicity (Non-Hispanic 90.3%, Hispanic 4.3%, Unknown 5.4%). Each image is paired with a binary glaucoma diagnosis label and a clinical question (“Does the patient have glaucoma?”). The dataset is split into 7,000 training, 1,000 validation, and 2,000 test samples, maintaining demographic distributions across splits.
@@
Model. We fine-tune Qwen2.5-VL-7B-Instruct 15 (8.3B parameters) with LoRA (, , dropout=) applied to query, key, value, and output projection layers () across all 32 attention layers. This configuration yields 20M trainable parameters (0.24% of the full model), striking a balance between expressiveness and parameter efficiency for fairness optimization on our medical dataset ( training samples). We freeze the vision encoder to preserve pre-trained visual features while adapting the language model for medical diagnosis.